# Multimodal Pathway: Improve Transformers with Irrelevant Data from Other Modalities

### 25 Jan 2024 v1

@@ 1 Introduction
Transformers [51, 19, 50, 22] are widely adopted in various tasks across modalities, such as text classification [10], map construction [65], object detection [4], point cloud analysis [66], and audio spectrogram recognition [24]. Apart from numerous unimodal tasks, transformers are also effective on multimodal data, e.g., CLIP [43] uses image-text pairs to achieve superior performance in image recognition. Transformers’ success in multiple modalities demonstrates their abilities to universally establish sequence-to-sequence modeling, given the input sequences (i.e., tokens) which can be seen as the universal embeddings of data of multiple modalities [20, 4, 58, 66, 24, 55]. For brevity, we refer to such ability as the universal modeling ability.
@@
Figure 2: (Left) Framework of Multimodal Pathway Transformer (M2PT). We use point cloud and image modalities as an example. Common practices with transformers follow the same pipeline: using 1) tokenizers to convert the input data to sequences, 2) transformer blocks to process the sequences, and 3) heads to decode the sequences. We upgrade the sequence-to-sequence modeling by establishing pathways between the components of different modalities so processing the tokens of a specific modality can utilize the transformer blocks trained with another modality. (Middle) Conceptual design of M2PT, where the pathways are implemented by letting a linear layer (including the Query/Key/Value/projection layers in the attention block and those in the FFN block) in the target model cooperate with its counterpart in the auxiliary model. (Right) Cross-Modal Re-parameterization efficiently realizes M2PT by re-parameterizing the weights of the target model with those of the auxiliary model, introduceing marginal training costs and completely no inference costs.
@@ 2 Related Work
Multimodal pretraining. Existing multimodal learning methods require paired [54, 56, 53, 55] or interleaved data [1]. In either case, the data samples of different modalities are well-aligned (i.e., strongly related). A recent study highlighted a main trend in the literature - existing multimodal pretraining methods are overly dependent on the well-aligned multimodal sample pairs/tuples [60]. For instance, VideoBERT [46] and CBT [45] utilize well-aligned video and speech data; many models are pretrained with large-scale image-text data, e.g., VisualBERT [34], VL-BERT [44], ViLBERT [39], LXMERT [48], UNITER [8], LLaVa [37], OneLLM [27], EVA-CLIP [47], GVT [52], VL-GPT [69].
Nowadays, using the weakly-aligned or unpaired/unaligned multimodal data as the pretraining corpora remains understudied [60]. This work represents an early exploration in this direction, which serves to fill this gap in the field.
Structural Re-parameterization is a methodology that constructs extra structures (e.g., convolutional layers) during training and converts the trained structures via transforming the parameters [12, 14, 13, 16, 11, 3, 15, 17, 18]. A primary drawback of Structural Re-parameterization is that the constructed layers must participate in the computations with the inputs, resulting in significant extra training costs.
@@ 3.1 Architectural Design
Image tokenizer. We represent an image by $\bm{x}_{I}\in\mathbb{R}^{H\times W\times C}$, where $(H,W)$ specifies the image’s resolution, and $C$ is the number of channels. Following ViT [20], we use an $S$$\times$$S$ patch embedding layer, which is implemented with an $S$$\times$$S$ convolutional layer with a stride of $S$, to project images into $D$-dimensional tokens. That is
@@ 4.1 Setup
Datasets. For image recognition, we evaluate the models’ performance on three representative image datasets. 1) ImageNet-1K [9] is the most widely adopted benchmark for visual perception tasks, which contains nearly 1.3 million images of 1000 categories. 2) MSCOCO 2017 [36] is a common benchmark for object detection. M2PT is trained on the train set and evaluated on the val set with Mask RCNN [28]. 3) ADE-20K [67] is used for semantic segmentation experiments with UperNet [57] and we adopt the single-scale evaluation setting. For point cloud, we evaluate the performance of M2PT on ShapeNetPart [61], which contains 16,880 models and 16 categories. For audio recognition, following AudioMAE [31], we utilize the AudioSet-2k [23] dataset. For video, we experiment on the action recognition dataset, Kinetics-400 [32], which contains 240k training videos and 20k validation videos from 400 classes.
Table 1: Experimental results on image recognition tasks. On ImageNet, we report the results with the linear layers in transformer blocks finetuned (tune acc) or fixed (fix acc). ***: results are reported by running the official code. The architecture of every model is ViT-B. The relative improvements over the baselines are shown in green.
@@
Experimental details. For a pair of target modality and auxiliary modality, we obtain the auxiliary model by self-supervised training on a dataset of the auxiliary modality. Specifically, the auxiliary image model is pretrained with MAE [30] on ImageNet-1K [9], the auxiliary point cloud model is pretrained with Point-MAE [41] on ShapeNet [6], the auxiliary audio model is pretrained with AudioMAE [31] on AudioSet-2M [23], the auxiliary video model is pretrained with VideoMAE [49] on Kinetics-700 [32]. For fairness and reproducibility, we use their official code for pretraining. We do not use supervised pretraining because we would like to eliminate the effects of labels in the pretraining datasets so that we can ensure the irrelevance of the data samples, considering that the labels of two datasets may somehow overlap. In terms of the initialization of the target model, we adopt two settings. 1) The target model (i.e., the parameters denoted by $\bm{W}$ in Eq. 9) is initialized with the aforementioned weights pretrained with the self-supervised methods on the target modality. We finetune the M2PT model with the default finetuning configurations described by the corresponding pretraining methods. The baseline model is also initialized with the pretrained weights and fine-tuned with identical configurations so that this setting is referred to as the pretrained setting for brevity. 2) The target model is randomly initialized as usual, and we use the widely adopted training configurations to train the M2PT model. The baseline model is trained from scratch with identical configurations for fair comparisons so that the setting is referred to as the from-scratch setting for brevity. In other words, the M2PT and baseline model both have no weights pretrained on the target modality under this setting.
Metrics. We report the performance of M2PT models on various datasets, including top-1 accuracy for ImageNet-1K, AudioSet, Kinetics-400, mIoU for ADE20K, ShapeNetPart and PartNet, and box/mask AP for MS COCO. To fairly assess the performance improvements over the baselines in multiple metrics, we report the relative percentage of improvement in Table 1234. For example, compared to the MAE-pretrained model, the ADE20K mIoU improves from 46.1 to 47.9 so that the relative improvement is $(47.9-46.1)/(46.1)=3.9\%$ (Table 1).
@@ 4.2 Main Results
Video understanding. For the experiments on Kinetics-400, we adopt only the pretrained setting because it is not a common practice to train a model from scratch on a video dataset, which would deliver inferior performance. We use the Video-MAE-pretrained ViT to initialize the target weights. Naturally, the baseline should be the VideoMAE-pretrained model directly finetuned on Kinetics-400. Table 4 shows that compared with previous works including SlowFast [21], MViTv2 [35], TimeSFormer [2], and VideoMAE [49], M2PT outperforms by at least +0.8 top-1 accuracy (82.3 vs. 81.5), which reveals that the temporal awareness for video understanding can also be enhanced with irrelevant data from other modalities.
@@ 4.3 Ablation Studies
Table 5: Ablation studies on design choices of M2PT including the layers to re-parameterize and configurations of Cross-Modal Scale $\lambda$. The target dataset is ImageNet-1K and the auxiliary modality is point cloud.
Components Cross-Modal Scale Top-1 accuracy (%)
Attn QKV Attn Proj FFN 1st FFN 2nd Init. Trainable
0 83.4
0 83.6
0 83.6
0 83.7
0 83.9
$10^{-2}$ 83.5
$10^{-2}$ 83.6
$10^{-4}$ 83.6
$10^{-6}$ 83.7
@@ 4.4.1 On the Modality-Complementary Knowledge
The observed improvements on multiple modalities have shown that the auxiliary transformer has learned some knowledge that is able to transfer to the target modality. We continue to investigate the properties of such modality-complementary knowledge through two groups of experiments (Table 6).
1) We investigate if such knowledge is related to the ability to generally process hierarchical representations. Abstraction hierarchy exists in multiple modalities with concepts ranging from low-level to high-level, which may explain the transferability of the learned knowledge. For example, in the image and point cloud modalities, this hierarchy may include textures (in images) or individual points (in point clouds), object parts, and whole objects. Considering that the conceptual level a transformer block works on is determined by its depth, we design an experiment by reverting the order of the auxiliary weights. Specifically, the counterpart of the first target block should be the first auxiliary block, whose weights are connected via Cross-Modal Re-parameterization, which is obvious. Under the reverse-order setting, since the transformer has 12 blocks, we let the $i$-th block connect with the $(13-i)$-th block so that the target-auxiliary correspondence is interrupted. We observe that doing so decreases the accuracy to 83.61%, which is 0.32% lower than the normal M2PT. In summary, we observe that modality-complementary knowledge in the auxiliary transformer can transfer to another modality but can be harmed if the low-to-high correspondence is interrupted, suggesting that such knowledge may help understand general hierarchical concepts regardless of the modality.
2) We investigate if the improvements are merely due to more trainable parameters or a better initialization by verifying if a better pretraining process brings the aforementioned knowledge of higher quality. We experiment using not well-trained weights as the auxiliary weights. Specifically, the default auxiliary weights are obtained through a 300-epoch self-supervised pretraining process on point cloud data, but we alternatively use the checkpoints saved at the 20th and 220th epoch, respectively, as the auxiliary weights. Not surprisingly, we observe that the performance degrades to 83.55% and 83.69%, respectively, which is still higher than the baseline. This phenomenon suggests that the improvements brought by the auxiliary weights cannot be simply explained that the weights trained on another modality merely offer an initialization hardly better than the random initialization or the model merely benefits from a larger number of trainable parameters (if so, training the checkpoint at the 20th epoch to 300 epochs would not bring observable eventual improvements on the target modality).
@@ 4.4.2 Discussion on the Data Scale
1) From small-scale data to large-scale data. Previous works such as Image2Point [59] and Point-CLIP [63] follow a common consensus that the modality owning a larger data scale could be utilized to benefit the other modality owning a smaller one. Therefore, Image2Point introduces image-pretrained models to data-insufficient 3D perception tasks. Differently, M2PT sets up a brand new methodology and breaks the former consensus - we discover that even though the data scale of point clouds is limited, such data still brings out impressive improvements to the image, video, and audio perception tasks. Impressively, the pretraining data of the latter modalities is larger in magnitude than that of the point cloud, but the point cloud data makes a difference. 2) From large-scale data to small-scale data. On the other hand, the effectiveness of M2PT highlights that for 3D vision research and other areas that lack large-scale data for pretraining, M2PT introduces a promising direction to leverage irrelevant large-scale data from other modalities.
@@ 5 Conclusion and Limitation
This paper explores the feasibility and advantages of improving a transformer’s performance on a specific modality with irrelevant data from other modalities. We propose a general framework named Multimodal Pathway and a concrete inference-cost-free implementation named Cross-Modal Re-parameterization. Multimodal Pathway represents an early exploration in this direction, which offers a novel perspective. We realize significant and consistent improvements on four representative modalities, demonstrating the potential of our method as a promising approach.
The primary limitation is that the theory behind the improvements remains to be revealed. Apart from empirical explanations, we believe further investigations (e.g., a mathematically provable bound) will be useful, which may require a deeper understanding of the black box of deep neural networks.

### 18 Mar 2024 v2

@@ 1 Introduction
Transformers [37, 12, 36, 14] are widely adopted in various tasks across modalities, such as text classification [8], object detection [3], point cloud analysis [47], and audio spectrogram recognition [16]. Apart from numerous unimodal tasks, transformers are also effective on multimodal data, e.g., CLIP [32] uses image-text pairs to achieve superior performance in image recognition. Transformers’ success in multiple modalities demonstrates their abilities to universally establish sequence-to-sequence modeling, given the input sequences (i.e., tokens) which can be seen as the universal embeddings of data of multiple modalities [12, 3, 47, 16, 46]. For brevity, we refer to such ability as the universal modeling ability.
@@
Figure 2: (Left) Framework of Multimodal Pathway Transformer (M2PT). We use point cloud and image modalities as an example. Common practices with transformers follow the same pipeline: using 1) tokenizers to convert the input data to sequences, 2) transformer blocks to process the sequences, and 3) heads to decode the sequences. We upgrade the sequence-to-sequence modeling by establishing pathways between the components of different modalities so processing the tokens of a specific modality can utilize the transformer blocks trained with another modality. (Middle) Conceptual design of M2PT, where the pathways are implemented by letting a linear layer (including the Query/Key/Value/projection layers in the attention block and those in the FFN block) in the target model cooperate with its counterpart in the auxiliary model. (Right) Cross-Modal Re-parameterization efficiently realizes M2PT by re-parameterizing the weights of the target model with those of the auxiliary model, introducing marginal training costs and completely no inference costs.
@@ 2 Related Work
Multimodal pretraining. Existing multimodal learning methods require paired [39, 40, 50, 19] or interleaved data [1]. In either case, the data samples of different modalities are well-aligned (i.e., strongly related). A recent study highlighted a main trend in the literature - existing multimodal pretraining methods are overly dependent on the well-aligned multimodal sample pairs/tuples [43]. For instance, VideoBERT [34] and CBT [33] utilize well-aligned video and speech data;
Nowadays, using the weakly-aligned or unpaired/unaligned multimodal data as the pretraining corpora remains understudied [43]. This work represents an early exploration in this direction, which serves to fill this gap in the field and contributes to multimodal calibration [38].
Structural Re-parameterization is a methodology that constructs extra structures (e.g., convolutional layers) during training and converts the trained structures via transforming the parameters [9, 10, 11]. A primary drawback of Structural Re-parameterization is that the constructed layers must participate in the computations with the inputs, resulting in significant extra training costs.
@@ 3.1 Architectural Design
Image tokenizer. We represent an image by $\bm{x}_{I}\in\mathbb{R}^{H\times W\times C}$, where $(H,W)$ specifies the image’s resolution, and $C$ is the number of channels. With an image patch of $(S,S)$, we obtain:
@@ 4.1 Setup
Datasets. For image recognition, we evaluate the models’ performance on three representative image datasets. 1) ImageNet-1K [7] contains nearly 1.3 million images of 1000 categories. 2) MSCOCO 2017 [27] is a common benchmark for object detection. M2PT is trained on the train set and evaluated on the val set with Mask RCNN [20]. 3) ADE-20K [48] is used for semantic segmentation experiments with UperNet [41] and we adopt the single-scale evaluation setting. For point cloud, we evaluate the performance of M2PT on ShapeNetPart [44], which contains 16,880 models and 16 categories. For audio recognition, following AudioMAE [23], we utilize the AudioSet-2k [15] dataset. For video, we experiment on the action recognition dataset, Kinetics-400 [24], which contains 240k training videos and 20k validation videos from 400 classes.
Table 1: Experimental results on image recognition tasks. On ImageNet, we report the results with the linear layers in transformer blocks finetuned (tune acc) or fixed (fix acc). The architecture of every model is ViT-B. The relative improvements over the baselines are shown in green. * The standard error of M2PT on image recognition tasks is 0.04.
@@
Experimental details. For a pair of target modality and auxiliary modality, we obtain the auxiliary model by self-supervised training on a dataset of the auxiliary modality. Specifically, the auxiliary image model is pretrained with MAE [22] on ImageNet-1K [7], the auxiliary point cloud model is pretrained with Point-MAE [30] on ShapeNet [5], the auxiliary audio model is pretrained with AudioMAE [23] on AudioSet-2M [15], the auxiliary video model is pretrained with VideoMAE [35] on Kinetics-700 [24]. We do not use supervised pretraining because we would like to eliminate the effects of labels in the pretraining datasets so that we can ensure the irrelevance of the data samples, considering that the labels of two datasets may somehow overlap. In terms of the initialization of the target model, we adopt two settings. 1) The target model (i.e., the parameters denoted by $\bm{W}$ in Eq. 9) is initialized with the aforementioned weights pretrained with the self-supervised methods on the target modality. We finetune the M2PT model with the default finetuning configurations described by the corresponding pretraining methods. The baseline model is also initialized with the pretrained weights and fine-tuned with identical configurations so that this setting is referred to as the pretrained setting for brevity. 2) The target model is initialized as usual, and we use the widely adopted training configurations to train the M2PT model. The baseline model is trained from scratch with identical configurations for fair comparisons so that the setting is referred to as the from-scratch setting for brevity.
Metrics. We report the performance of M2PT models on various datasets, including top-1 accuracy for ImageNet-1K, AudioSet, Kinetics-400, mIoU for ADE20K, ShapeNetPart and PartNet, and box/mask AP for MS COCO. To fairly assess the performance improvements over the baselines in multiple metrics, we also report the relative percentage of improvement in Table 123, and 4.2.
@@ 4.2 Main Results
Video understanding. For the experiments on Kinetics-400, we adopt only the pretrained setting because it is not a common practice to train a model from scratch on a video dataset, which would deliver inferior performance. We use the Video-MAE-pretrained ViT to initialize the target weights. Naturally, the baseline should be the VideoMAE-pretrained model directly finetuned on Kinetics-400. Table 4.2 shows that compared with previous works including SlowFast [13], MViTv2 [26], TimeSFormer [2], and VideoMAE [35], M2PT outperforms by at least +0.8 top-1 accuracy (82.3 vs. 81.5), which reveals that the temporal awareness for video understanding can also be enhanced with irrelevant data from other modalities.
@@ 4.3 Ablation Studies
Table 5: Ablation studies on design choices of M2PT including the layers to re-parameterize and configurations of Cross-Modal Scale $\lambda$. We use the point cloud and video as auxiliary modalities for image and 3D evaluation. The first row reports the results of direct tuning.
Multimodal Pathway Components Cross-Modal Scale ImageNet ShapeNetPart PartNet
Attn QKV Attn Proj FFN 1st FFN 2nd Init. Trainable (%) (%) (%)
- - 83.3 84.2/86.1 47.4
0 83.4 84.6/86.5 48.3
0 83.6 84.8/87.1 48.2
0 83.6 84.9/87.0 48.4
0 83.7 85.2/87.2 48.3
0 83.9 85.6/87.5 50.1
$10^{-2}$ 83.5 84.6/86.3 48.2
$10^{-2}$ 83.6 84.3/86.2 48.0
$10^{-4}$ 83.6 84.7/86.2 48.1
$10^{-6}$ 83.7 84.7/86.4 48.2
@@
Table 7: Training efficiency of Multimodal Pathway.
Model Train Time Train Param. Inference Time Inference Param.
MAE 16.95 Hours 86.3M 11.64 ms 86.3M
M2PT 22.84 Hours 172.6M 11.64ms 86.3M
@@ 4.4.1 On the Modality-Complementary Knowledge
The observed improvements on multiple modalities have shown that the auxiliary transformer has learned some knowledge that can be transferred to the target modality. We continue to investigate the properties of such modality-complementary knowledge through two groups of experiments (Table 6).
1) Modality-complementary knowledge & Abstraction Hierarchy. Vision Transformers excel in general hierarchical representations by stacking blocks [12]. For example, in the image and point cloud modalities, this hierarchy may include textures (in images) or individual points (in point clouds), object parts, and whole objects. In Table 6, we construct the multimodal pathway by connecting transformer blocks of different depths. Specifically, the counterpart of the first target block should be the first auxiliary block. Under the reverse-order setting, we observe that doing so decreases the accuracy to 83.61%, which is 0.32% lower than the normal M2PT. We observe that modality-complementary knowledge in the auxiliary transformer can transfer to another modality but can be harmed if the low-to-high correspondence is interrupted, suggesting that modality-complementary knowledge reinforces hierarchical representations of the transformer architecture.
2) More trainable parameters? Just better initialization? For this part, we use insufficiently pretrained auxiliary weights. Specifically, the default auxiliary weights are pretrained for 300 epochs with mask modeling on point cloud data, but we alternatively use the checkpoints saved at the 20th and 220th epoch, respectively, as the auxiliary weights. Not surprisingly, we observe that the performance degrades to 83.55% and 83.69%, respectively, which is still higher than the baseline. This phenomenon suggests that the improvements brought by the auxiliary weights cannot be explained as better initialization, because after pretraining the auxiliary model from 20 to 300 epochs, the accuracy increases from 83.5 to 83.9. If improvements were due to initialization, the results of pretraining 20 epochs should be close to random initialization (83.5 v.s. 81.9).
@@ 4.4.2 Discussion on the Data Scale
1) From small-scale data to large-scale data. Previous works such as Image2Point [42] introduces image-pretrained models to data-insufficient 3D perception tasks. Differently, M2PT sets up a brand new methodology and breaks the former consensus - we discover that even though the data scale of point clouds is limited, such data still brings out impressive improvements to the image, video, and audio perception tasks. Impressively, the pretraining data of the latter modalities is larger in magnitude than that of the point cloud, but the point cloud data makes a difference. 2) From large-scale data to small-scale data. On the other hand, the effectiveness of M2PT highlights that for 3D vision research and other areas that lack large-scale data for pretraining, M2PT introduces a promising direction to leverage irrelevant large-scale data from other modalities.
@@ 5 Conclusion and Limitation
This paper explores the feasibility and advantages of improving a transformer’s performance on a specific modality with irrelevant data from other modalities. We propose the Multimodal Pathway and a concrete implementation of no additional inference cost named Cross-Modal Re-parameterization. It represents an early exploration in this direction, which offers a novel perspective. We realize significant and consistent improvements on four representative modalities, demonstrating the potential of our method as a promising approach. In the future, we will explore to construct multimodal pathways among CNNs and cross-architecture. The primary limitation is that the theory behind the improvements remains to be revealed. Apart from empirical explanations, we believe further investigations (e.g., a mathematically provable bound) will be useful.
Acknowledgements. This work is partially supported by the National Natural Science Foundation of China (Grant No. 8326014).